OpenStudy (montesanna99):

Verify the identity cos(x+pi/2)= -sinx

OpenStudy (mathmale):

Use the sum formula for the cosine function:\[\cos (a+b) = \cos a \cos b - \sin a \sin b\] Look up cos pi/2 and sin pi/2.

OpenStudy (montesanna99):

im sorry im a little confused

OpenStudy (mathmale):

The identity I gave you is always true. Substitute x for a and pi/2 for b. Write out the identity after having done this. Use the fact that sin pi/2=1 to simplify your equation. Also use the fact that cos pi/2=0 to simplify it.

OpenStudy (kevin):

Just to make you're not confused : cos(x + pi/2) = cos x cos pi/2 - sin x sin pi/2 = cos x cos 90 - sin x sin 90 = cos x (0) - sin x (1) = 0 - sin x = -sin x

OpenStudy (montesanna99):

Thank you, everyone